1. The average number of milligrams (mg) of cholesterol in a cup of a certain brand of ice cream is





660 mg, the standard deviation is 35 mg. Assume the variable is normally distributed. a. If a cup of





ice cream is selected, what is the probability that the cholesterol content will be more than 670 mg?

1
Expert's answer
2022-04-11T16:34:07-0400

"P(X>670)=P(Z>\frac{670-660}{35})=P(Z>0.29)=1-P(Z<0.29)=0.3859."